From a square piece of cardboard, a box with an open top is made by cutting equal squares from the corners, and folding up the sides. If the height of the box is 2 inches and the volume is 72 cubic inches, find the length of one side of the cardboard.
Answer by oberobic(2304)

L = length of the cardboard
The cardboard is square
L^2 = area of the cardboard
.
L -4 = length of the sides of the square box made from the cardboard.
It is L-4 because you have to cut a 2x2 square from each corner to make the box.
.
Volume of the box = 72 cubic in.
.
V = (L-4)*(L-4)*2
V = 72
(L-4)^2 = 36
L-4 = 6
L = 10
.
So, the cardboard square originally had sides = 10 inches.
.
Done.